lalizing In consumer electronics. A condensed income statement and balance sheet for the Bargain Deal, Inc., is a leading retailer spec fiscal year ended January 30, 2016, are shown below. 1, Inc Balance Sheet At January 30, 2016 Assets Current assetS1 Cash and cash equivalents Short-term investments Accounts receivable, net Merchandise inventories Other current assets 2,136 1,321 1,242 5,067 424 10,190 3, 713 Total current assetB Long-term assets Total assets $13,903 Liabilities and Shareholder Equity Current liabilities Accounts payable 5,250 4,075 9,325 2,248 2,330 $13,903 other current liabilities Total current liabilities Long-term liabilitien Shareholders equity Total liabilities and shareholdera equity Bargain Deal, Inc Income Statement For the Year Ended January 30, 2016 ($ in milliona) $39,608 38,169 1,439 81 1,358 743 s 615 Revenuen Costa and expenses Operating income Other income (expense)* Income before income taxes Income tax expense Net income Includes $224 of interest expense
Bargain Deal, Inc Income Statement For the Year Ended January 30, 2016 ($ in millions r Inc Revenues $39,608 38,169 1,439 81 1,358 743 Costs and expenses Operating income Income before income taxes Net income Other income (expense)* Income tax expense $ 615 Includes $224 of interest expense. Required: 1-a. C 1-b. C alculate the current ratio for Bargain Deal for its fiscal year ended January 30, 2016. alculate the acid-test ratio for Bargain Deal for its fiscal year ended January 30, 2016. 1c. Calculate the debt to equity ratio for Bargain Deal for its fiscal year ended January 30, 2016. 1-d. Calcu late the times interest earned ratio for Bargain Deal for its fiscal year ended January 30, 2016. (For all requirements, round your answers to 2 decimal places.) 1-a. Current ratio 1-b. Acid-test ratio 1-c. Debt to equity ratio 1-d. Times interest earned ratio times
Answer #1
1-a) Current ratio = Current assets/Current liabilities = 10190/9325 = 1.09
1-b) Acid-test ratio = (Cash & cash equivalents+Short term investments+Accounts receivable)/Current liabilities = (2136+1321+1242)/9325 = 0.50
1-c) Debt to equity ratio = Total debt/Shareholders' equity = (9325+2248)/2330 = 4.97
1-d) Times interest earned ratio = Operating income/Interest expense = 1439/224 = 6.42 times